OpenStudy (anonymous):

At what temperature would the velocities of CO2 and Neon be equal?

OpenStudy (jfraser):

no single temperature (except absolute zero) will have 2 gasses of different masses have the same velocity. If the two gasses are at the same temp, then their average kinetic energies are the same. However, since \[KE = \frac{1}{2}mv^2\] then\[KE_{CO_2} = \frac {1}{2}m_{CO_2}v_{CO_2}^2 = \frac{1}{2}m_{Ne}v_{Ne}^2 = KE_{Ne}\]Since the masses are different, the velocities must be different in order to maintin the equality of their KEs. If you had the temp and speed of one gas, you could ask what would the temp of the other have to be to match speeds, but the way it's stated, the only logical answer is either no temp, or absolute zero